find the minimal polynomial of the linear operator t : R³ "- R³" define by t (x,y,z) =(x+2y+3z, 4y+5z,6 z).is t




1
Expert's answer
2022-02-07T13:58:06-0500

"\begin{pmatrix}
 X+2y+3z \\
 4y+5z\\
6z
\end{pmatrix}" "=x\begin{pmatrix}
 1\\
 0\\
0
\end{pmatrix}+y\begin{pmatrix}
 2\\
 4 \\
0
\end{pmatrix}+z \begin{pmatrix}
 3 \\
 5 \\
6
\end{pmatrix}"



Transforming matrix

"A=\begin{pmatrix}
 1&2&3 \\
 0&4&5\\
0&0&6
\end{pmatrix}"


Characteristic equation of A is

"\>\>\begin{vmatrix}
 A-I\lambda \\
 
\end{vmatrix}=\begin{vmatrix}
 1-\lambda&&2&&3 \\
 0&&4-\lambda&&5 \\
0&&0&&6-\lambda
\end{vmatrix}=0"


Expanding

"(1-\lambda)[(4-\lambda)(6-\lambda)-0]+2(0)+3(0)=0"


"\implies(1-\lambda)(4-\lambda)(6-\lambda)=0"


The distincts roots of A are also roots of the minimum polynomial "g(x)"


"\therefore\>g(x)=(x-1)(x-4)(x-6)"